For angular simple harmonic motion or physical pendulum torque is calculated about which point or line.( in reference to hc verma part 1 chapter 12 question 26)

Grade:Upto college level

2 Answers

aku -- kumar
13 years ago

in this kind of problems consider oscillation of centre of mass only that will make ur problem easier to solve

 

let the pendulum be at some Φ angle with vertical axis,  draw f.b.d ur self,

 

balance the driving force with mg,i.e; F=mg sinΦ 

so a= F/m=g sinΦ if Φis very small then sinΦ ≈ Φ, 

a = gΦ =g(x/l)     » a  proportional to x

 

then use time period formulae, t=2∏ √x/a .

Aravind Pazhayam
12 years ago

In a physical pendulum the pendulum is hinged about a point and it rotates about that point.........however in this case this point cant be CM ...so the torque has to be take about a point where it does not pass through the line of action......in case of a mathematical pendulum it moves about the pivoted point.................for deep concepts refer RESNICK AND HALLIDAY......pardon me if in any way im wrong.....